A fly appears to be upside down and twice its actual size when viewed through a converging lens with a focal length of 5.0 cm. How far is the fly from the lens?
Answer #1

Since the image is inverted

M = -2 = -s'/s

so s' = 2s

s' ,s = image , object distance

f = 5 cm

1/s + 1/s' = 1/f

1/s + 1/(2s) = 1/5

1/s(1 + 1/2)

=1/s(1.5)

= 1.5/s = 1/5

s = 1.5* 5 = 7.5 cm
The fly is 7.5 cm in front of the lens.

I hope help you !!